# Characterizing LLM Abstention Behavior in Science QA with Context Perturbations

### v1

@@ Abstract
The correct model response in the face of uncertainty is to abstain from answering a question so as not to mislead the user. In this work, we study the ability of LLMs to abstain from answering context-dependent science questions when provided insufficient or incorrect context. We probe model sensitivity in several settings: removing gold context, replacing gold context with irrelevant context, and providing additional context beyond what is given. In experiments on four QA datasets with four LLMs, we show that performance varies greatly across models, across the type of context provided, and also by question type; in particular, many LLMs seem unable to abstain from answering boolean questions using standard QA prompts. Our analysis also highlights the unexpected impact of abstention performance on QA task accuracy. Counter-intuitively, in some settings, replacing gold context with irrelevant context or adding irrelevant context to gold context can improve abstention performance in a way that results in improvements in task performance. Our results imply that changes are needed in QA dataset design and evaluation to more effectively assess the correctness and downstream impacts of model abstention. [^1] Code will be publicly available at https://github.com/bbwen/llm_scienceqa.
@@ 1 Introduction
Question-answering (QA) in scientific settings is typically defined as a context-dependent task, where models answer questions based on provided context or relevant context it identifies. When the provided or retrieved context is itself unreliable or inconsistent, however, the correct model response should be to abstain from answering. Prior work has studied the abstention capabilities of LLMs Yin et al. (2023); Amayuelas et al. (2023) and proposed approaches to improve abstention when presented with insufficient context Zhou et al. (2023); Slobodkin et al. (2023), but these approaches focus on general domain question-answering (e.g. SQuAD2) against a narrow range of models (e.g. ChatGPT). As we will show in our experiments (§5.2), general domain settings produce highly divergent results from scientific QA settings, and different model architectures also exhibit differing abilities to abstain. Further, as new LLMs continue to be developed and released, we need an extensible way to measure their ability to abstain from answering questions when provided context is irrelevant or uncertain.
To address these challenges, we introduce a framework to assess LLM abstention in science question-answering by removing, replacing, and augmenting provided contexts to control the answerability of questions. Using this framework, we probe abstention performance of four LLMs (LLama2, Vicuna, Flan-T5, and GPT3.5) on one general domain QA dataset (SQuAD2) and three scientific QA datasets (PubmedQA, BioASQ, QASPER). Several of these datasets include unanswerable questions by design, allowing us to analyze interactions between question answerability and noisy context. Our work aims to answer the following: (i) How well do LLMs abstain from answering questions when the correct context is not provided? (ii) How do context perturbations impact task performance and abstention performance? And (iii) How does question type impact task performance and abstention performance? We summarize our contributions below:
@@
  • Using this framework, we investigate the task and abstention performance of four LLMs on four QA datasets, ranging from general-domain factoid QA to context-sensitive, document-based science QA (§5). Our results show that no model consistently abstains in all settings where abstention is expected (unanswerable questions and no context/random context settings), though some models demonstrate a stronger ability to abstain for more context-dependent QA tasks, and instruction-tuned models are generally better at following abstention instructions.
@@ 2 Related Work
The few-shot capability of LLMs have been applied successfully to knowledge-intensive tasks like question answering (Wei et al., 2021; Chowdhery et al., 2022; Nori et al., 2023). Prompting strategies such as chain-of-thought (Wei et al., 2022), least-to-most (Zhou et al., 2022), and others (Kojima et al., 2022; Wang et al., 2022) significantly improve LLMs’ zero or few-shot abilities on diverse QA benchmarks in the general Rajpurkar et al. (2018); Yang et al. (2015) and scientific Taylor et al. (2022); Pereira et al. (2023) domains. Pereira et al. (2023) adopt retrieve-than-read on the QASPER dataset Dasigi et al. (2021), showing that current retrievers are the main bottleneck and readers (LLMs) are already performing at the human level. Inspired by this finding, we study how irrelevant or incorrect contexts (mimicking retrieval errors) can impact LLM performance on context-intensive QA tasks.
LLMs and Abstention in QA
Several works have shown that LLMs become easier to steer with natural language prompts either as they become larger Mishra et al. (2022a) or as they are exposed to more instruction tuning Mishra et al. (2022b); Chung et al. (2022). Liao et al. (2022) introduced a prompt tuning strategy to improve performance on unanswerable questions, by mapping questions onto their specific templates. Other work tried to direct models toward abstention through data augmentation Zhu et al. (2019). Asai and Choi (2021) provided an in-depth analysis of LMs’ abstention ability, identifying paragraph selection and answerability prediction as two areas for improvement. Recent work introduced new datasets to study whether LLMs know what they do not know Yin et al. (2023); Amayuelas et al. (2023), while Slobodkin et al. (2023) demonstrated that differences in LLM hidden states identify the boundary between known and unknown. In this work, we systematically characterize LLM abstention capabilities on context-dependent QA tasks.
@@
Prior work studying input perturbations for NLP tasks include approaches such as model-agnostic input transformations Liang et al. (2022); Ravichander et al. (2022); Giorgi et al. (2022) and adversarial example generation Jia and Liang (2017); Wang et al. (2021). Liang et al. (2022) use semantics-preserving and semantics-altering perturbations in their robustness evaluation of LLMs. Pretrained language models can be negatively impacted by irrelevant context (Chowdhery et al., 2022; Liang et al., 2022), e.g., Shi et al. (2023) injected irrelevant numerical context for the MathQA dataset, after which ChatGPT performance dramatically decreased. However, Liang et al. (2022) evaluated T5 Raffel et al. (2020) and PaLM Chowdhery et al. (2022), demonstrating that finetuning these models with counterfactual and irrelevant contexts can improve model robustness to noisy context. In our framework, we leverage context perturbations to investigate the LLM abstention behavior for science QA, finding occasional counter-intuitive interactions between abstention and task performance.
@@ 4.3 Models
We conduct experiments using LLamaV2-13b-chat (LLama2), Vicuna1.5-13b-chat (Vicuna), Flan-T5-XL (Flan-T5), and gpt-3.5-turbo-0613 (GPT3.5). We select these models to have reasonable representation across the following attributes:
  • Closed API (GPT3.5) vs open weights (LLama2, Vicuna, and Flan-T5)
  • Encoder-decoder (Flan-T5) vs decoder-only (LLama2, Vicuna, and GPT3.5) architecture
@@
Post-processing model output LLMs usually produce long outputs, so we post-process their outputs to obtain structured predicted answers. For boolean questions, we extract the first words from the model output. If these words contain “yes,” “no,” or “unanswerable,” we map them to the corresponding classes.
@@ 4.4 Evaluation metrics
Table 3: Model zero-shot task performance using constrained prompts. SOTA indicates the best zero-shot performance of LLMs reported in previous papers (*=Flan-UL2 Slobodkin et al. (2023), **=Galactica Taylor et al. (2022)), or best performance from a pretrained LM ($\dagger$=UnifiedQA-large Dasigi et al. (2021)). Best performances bolded, second best underlined among baseline performance (a). Colors indicate positive or negative delta from baseline task performance with different context perturbations: (b) no context, (c) random context, and (d) noisy context. While task performance generally degrades with context perturbations, this is not the case for QASPER due to interactions between abstention and task performance (see Section 5.1).
@@ 5 Results
Baseline task performance in the zero-shot prompting setting is presented in Table 3(a), alongside previously reported SOTA zero-shot performance of LLMs on each dataset. Surprisingly, Flan-T5 achieves the best performance on SQuAD2, PubmedQA and QASPER, LLama2 performs best on BioASQ. GPT3.5 achieves the second best performance (very close to the best one) on PubmedQA, BioASQ and QASPER. The task performance of tested models are close to and sometimes exceed the reported SOTA on each dataset. Though in this work, we focus on assessing the change to task performance under different context settings rather than maximizing model performance.
@@ 5.1 Impact of context perturbations
In Table 4(c), we observe that all models are much more likely to abstain from answering when given random context except on the BioASQ dataset. For SQuAD2 and PubmedQA, random context improves most models’ abstention performance to close to 1. Specifically, random context improves Flan-T5’s ability to abstain from answering on PubmedQA while the no context setting fails to do so. For QASPER, the abstention performance on answerable questions in the random context setting is much higher compared to no context, but is still far from 1. Oddly, for boolean questions in BioASQ, random context biases the model towards answering “no” rather than “unanswerable”, and we observe no changes in abstention rate (further analysis in Section 6).
@@
While noisy context has mixed effects on abstention (Table 4(d)), this perturbation does not always translate to negative impacts on task performance. In Table 3(d), task performance increases on QASPER for LLama2, Vicuna, and Flan-T5. Since QASPER’s unanswerable questions have empty context, the baseline task performance for unanswerable questions is actually computed with no context; perturbing no context by adding noisy context therefore leads to a trade-off in task performance and abstention performance (some models abstain much more for unanswerable questions and task performance consequently improves).
@@ 5.2 Impact of question type
For extractive questions, Figure 2(a) and (d) show that model abstention varies significantly between given context and no context settings; under no context settings, models can achieve abstention performance close to 1. Abstractive answers show similar patterns. Notably, we find all models are reluctant to abstain from boolean questions for all datasets; this overconfidence is apparent even for QASPER, which is a highly document-specific QA dataset. Surprisingly, Flan-T5 and LLama2 demonstrate near zero abstention performance on boolean questions on any dataset, regardless of whether or not context is provided. [^2] We conduct prompt ablations in Appendix C, showing the lack of abstention on boolean questions is highly sensitive to prompt wording. Without a boolean instruction in the prompt, both models abstain well for boolean questions. Vicuna and GPT3.5 perform only slightly better on this front.
@@
In Figure 2, we examine the abstention performance on unanswerable questions (yellow background) and find that performance varies significantly across datasets and models. For SQuAD2, GPT3.5 performs consistently regardless of whether context is provided, indicating that GPT3.5 may be able to recognize the irrelevance of the context. Surprisingly, LLama2 and Vicuna abstain less when context is removed, which contradicts intuition. In SQuAD2, unanswerable questions are still related to the provided context; which may explain why when the confusing context is removed, these two models become more likely to answer. For PubmedQA, unanswerable questions have high-uncertainty context; Flan-T5 and LLama2 consistently refuse to abstain from answering these questions, which are boolean, while Vicuna follows the “unanswerable” instruction quite well. For QASPER, abstention performance is generally higher across all models compared to SQuAD2 and PubmedQA since the questions are more document-grounded.
@@
Table 5: % of answerable questions for which the model changes from answering to abstaining when context is removed. Best performances bolded, second best underlined.
5.3 Abstention performance of LLMs
@@
Table 5 shows the proportion of answers for which the model changed from answering to abstaining when context is removed. The expectation for context-dependent questions would be that a model would abstain from answering if no context is provided. GPT3.5 achieves the largest abstention changes on PubmedQA, BioASQ, and QASPER, indicating that it is responsive to context loss. Vicuna, as an instruction-tuned model, also demonstrates relatively strong abstention behavior in these cases, though to a lesser degree than GPT3.5 for all datasets. This effect is contrary to LLama2, which abstains at a similar rate regardless of whether context is provided.
@@
Flan-T5 abstains well when context is removed (Table 5). We observe large changes for SQuAD2 and QASPER (comparable or better than GPT3.5), although this behavior does not generalize at all to boolean questions—the change is 0.0 for PubmedQA and BioASQ, which consist only of boolean questions. Different prompting strategies may be necessary to enjoin certain models to abstain when answering boolean questions.
@@ 6 Error Analysis
We perform error analysis on the most consistently high performing model in our analysis, GPT3.5, and summarize main reasons why the model may fail to abstain. We sample 20 failure cases each from among answerable questions and unanswerable questions, for both gold context and no context settings. We present our findings by dataset.
@@ 7 Discussion & Conclusion
Our study investigates the impacts of context removal and perturbation on LLM performance on scientific QA. While lack of correct context should result in a model abstaining from answering a question, our results highlight that there are inconsistent patterns of model behavior based on model pretraining paradigms, question types, and the context-dependence of various QA datasets. For example, perturbing given context by substituting with random irrelevant context or adding noisy context would be expected to reduce task performance, but in some cases, the improvements in abstention that result can negate any reductions in task performance and potentially lead to gains (e.g., LLama2, Vicuna, and Flan-T5 on QASPER). Additionally, we find that abstention varies greatly by question type, with all models in our experiments struggling to abstain on boolean questions.
@@
Future work for enhancing models’ abstention ability could investigate the impact of (i) Different prompting strategies: since model abstention ability is sensitive to constrained prompts (such as boolean instructions) as shown in Section 5.2 and Appendix C, how to select good prompting strategies to produce the best trade-off between task performance and abstention performance remains an interesting problem; (ii) Alternate model architectures: smaller models such as Flan-T5-XL with encoder-decoder architectures performed comparably to larger decoder-only models such as ChatGPT. Further exploration of encoder-decoder architectures or introducing an auxiliary module to foster understanding of context may be helpful; (iii) Other context perturbations: we show initial results that providing noisy context can counter-intuitively improve task performance in some datasets due to interactions with unanswerable questions. How this interacts with retrieval errors that occur in open-domain QA is a direction that could be explored in future work.
@@
Our results have direct implications for dataset curators and model developers. Benchmark QA datasets with mixtures of unanswerable and answerable questions were designed to facilitate assessment of abstention ability, yet our experiments show conflation and a lack of clear assessment of either abstention ability or performance accuracy. In other words, while unanswerable questions were motivated by the need to measure abstention, when aggregated with other questions during task performance evaluation, this distinction is lost. Additionally, unanswerable questions in different datasets measure different phenomena, e.g., in SQuAD2, they measure model sensitivity to irrelevant context passages, in QASPER, they indicate that a question is unanswerable based on the given document, neither of which clearly map to the notion of abstaining under insufficient information. Given the importance of assessing model abstention capabilities, separating task and abstention assessment, coupled with changes in dataset construction, is needed to better align model performance on these tasks with human expectations.
@@
On the other hand, our results also hold implications for downstream builders/users of interactive systems who rely on LLMs as question-answering tools. We show that in some cases (e.g., boolean questions), LLMs exhibit little to no ability to abstain, even for highly context-dependent questions such as those in QASPER. System designers should be cautious of these limitations and when/how to infer model uncertainty and communicate this uncertainty to users. For communicating abstention capabilities to users, confidence scores or similar indicators that reflect model certainty, accompanying explanations, or guidance on how to interpret model answers can empower users to make more informed decisions.
@@
Due to the budget constraints, we could not conduct experiments with every single model, dataset, or context perturbation method from existing work, though our framework can be applied to other datasets and models. We will make our framework publicly available such that new models and datasets, as well as context perturbations could be added. Other factors affecting abstention such as in-domain pretraining and in-context learning were also not studied in this work and could be incorporated in the future.

### v2

@@ Abstract
The correct model response in the face of uncertainty is to abstain from answering a question so as not to mislead the user. In this work, we study the ability of LLMs to abstain from answering context-dependent science questions when provided insufficient or incorrect context. We probe model sensitivity in several settings: removing gold context, replacing gold context with irrelevant context, and providing additional context beyond what is given. In experiments on four QA datasets with six LLMs, we show that performance varies greatly across models, across the type of context provided, and also by question type; in particular, many LLMs seem unable to abstain from answering boolean questions using standard QA prompts. Our analysis also highlights the unexpected impact of abstention performance on QA task accuracy. Counter-intuitively, in some settings, replacing gold context with irrelevant context or adding irrelevant context to gold context can improve abstention performance in a way that results in improvements in task performance. Our results imply that changes are needed in QA dataset design and evaluation to more effectively assess the correctness and downstream impacts of model abstention. [^1] Code is publicly available at https://github.com/bbwen/llm_scienceqa.
@@ 1 Introduction
Question-answering (QA) in scientific settings is often defined as a context-dependent task, where models answer questions based on provided context or relevant context it identifies. When the provided or retrieved context is itself unreliable or inconsistent, however, the correct model response should be to abstain from answering. Prior work has studied the abstention capabilities of LLMs Yin et al. (2023); Amayuelas et al. (2024) and proposed approaches to improve abstention when presented with insufficient context Zhou et al. (2023); Slobodkin et al. (2023), but these approaches focus on general domain question-answering (e.g., SQuAD2) against a narrow range of models (e.g., ChatGPT). As we show in our experiments (§5.2), general domain settings produce highly divergent results from scientific QA settings, and different model architectures also exhibit differing abilities to abstain. Further, as new LLMs continue to be developed and released, we need an extensible way to measure their ability to abstain from answering questions when provided context is irrelevant or uncertain.
To address these challenges, we introduce a framework to assess LLM abstention in science question-answering by removing, replacing, and augmenting provided contexts to control the answerability of questions. Using this framework, we probe abstention performance of six LLMs (LLama2, Vicuna, Flan-T5, Mistral, GPT3.5, and GPT4o) on one general domain QA dataset (SQuAD2) and three scientific QA datasets (PubmedQA, BioASQ, QASPER). Several of these datasets include unanswerable questions by design, allowing us to analyze interactions between question answerability and noisy context. Our work aims to answer the following: (i) How well do LLMs abstain from answering questions when the correct context is not provided? (ii) How do context perturbations impact task performance and abstention performance? And (iii) How does question type impact task performance and abstention performance? We summarize our contributions below:
@@
  • Using this framework, we investigate the task and abstention performance of six LLMs on four QA datasets, ranging from general-domain factoid QA to context-sensitive, document-based science QA (§5). Our results show that no model consistently abstains in all settings where abstention is expected (unanswerable questions and no context/random context settings), though some models demonstrate a stronger ability to abstain for more context-dependent QA tasks, and instruction-tuned models are generally better at following abstention instructions.
@@ 2 Related Work
The few-shot capability of LLMs have been applied successfully to knowledge-intensive tasks like question answering (Wei et al., 2021; Chowdhery et al., 2024; Nori et al., 2023). Prompting strategies such as chain-of-thought (Wei et al., 2024), least-to-most (Zhou et al., 2022), and others (Kojima et al., 2024; Wang et al., 2022) significantly improve LLMs’ zero or few-shot abilities on diverse QA benchmarks in the general Rajpurkar et al. (2018); Yang et al. (2015) and scientific Taylor et al. (2022); Pereira et al. (2023) domains. Pereira et al. (2023) adopt retrieve-than-read on the QASPER dataset Dasigi et al. (2021), showing that current retrievers are the main bottleneck and readers (LLMs) are already performing at the human level. Inspired by this finding, we study how irrelevant or incorrect contexts (mimicking retrieval errors) can impact LLM performance on context-intensive QA tasks.
Abstention in LLMs
Liao et al. (2022) introduced a prompt-tuning strategy to enhance performance on unanswerable questions by mapping questions to specific templates. Beyond prompting strategies, Cole et al. (2023) demonstrated that sampling methods are more effective at modulating uncertainty compared to relying on the model’s likelihood. Other approaches, such as data augmentation, have been used to steer models toward abstention Zhu et al. (2019). Asai and Choi (2021) provided a detailed analysis of language models’ abstention capabilities, identifying paragraph selection and answerability prediction as key areas for improvement. Recent work has introduced new datasets to investigate whether LLMs are aware of their limitations Yin et al. (2023); Amayuelas et al. (2024), while Slobodkin et al. (2023) revealed that differences in LLM hidden states can delineate the boundary between known and unknown information. Abstention is also closely linked to out-of-distribution (OOD) detection and selective generation. OOD scores derived from conditional language models Ren et al. (2022) or task representations Chen et al. (2023) have been employed to abstain from generating low-quality outputs. In this work, we systematically characterize LLM abstention capabilities in context-dependent science question-answering tasks.
@@
Prior work studying input perturbations for NLP tasks include approaches such as model-agnostic input transformations Liang et al. (2022); Ravichander et al. (2022); Giorgi et al. (2022) and adversarial example generation Jia and Liang (2017); Wang et al. (2021). Liang et al. (2022) use semantics-preserving and semantics-altering perturbations in their robustness evaluation of LLMs. Pretrained language models can be negatively impacted by irrelevant context (Chowdhery et al., 2024; Liang et al., 2022), e.g., Shi et al. (2023) injected irrelevant numerical context for the MathQA dataset, after which ChatGPT performance dramatically decreased. However, Liang et al. (2022) evaluated T5 Raffel et al. (2020) and PaLM Chowdhery et al. (2024), demonstrating that finetuning these models with counterfactual and irrelevant contexts can improve model robustness to noisy context. Additionally, Liu et al. (2023) created a benchmark to systematically assess LLMs’ resilience against counterfactual knowledge derived from external texts, while Zou et al. (2024) proposed an attack method that introduces malicious texts into the knowledge base of a RAG system to manipulate the LLM into generating an attacker-specified response. In our framework, we leverage context perturbations to investigate the LLM abstention behavior for science QA, finding occasional counter-intuitive interactions between abstention and task performance.
@@ 4.3 Models
We conduct experiments using LLamaV2-13b-chat (LLama2), Vicuna1.5-13b-chat (Vicuna), Flan-T5-XL (Flan-T5), Mistral-7B-Instruct-v0.1 (Mistral), gpt-3.5-turbo-0613 (GPT3.5), and gpt-4o-2024-05-13 (GPT4o). We select these models to have reasonable representation across the following attributes:
  • Closed API (GPT3.5; GPT4o) vs open weights (LLama2, Vicuna, Flan-T5, and Mistral)
  • Encoder-decoder (Flan-T5) vs decoder-only (LLama2, Vicuna, Mistral, GPT3.5, and GPT4o) architecture
@@
Post-processing model output LLMs often generate lengthy responses, so we post-process to extract structured answers. For boolean questions, we take the first words of the output and map them to "yes," "no," or "unanswerable" based on their occurrence.
@@ 4.4 Evaluation metrics
GPT4o -0.6 +1.4 -0.8 -5.2
Table 3: Model zero-shot task performance using constrained prompts. SOTA indicates the best zero-shot performance of LLMs reported in previous papers (*=Flan-UL2 Slobodkin et al. (2023), **=Galactica Taylor et al. (2022)), or best performance from a pretrained LM ($\dagger$=UnifiedQA-large Dasigi et al. (2021)). (a) Among baseline model performance with the given context, the best performances are bolded, second best underlined. Colors indicate positive or negative delta from baseline task performance with different context perturbations: (b) no context, (c) random context, and (d) noisy context. While task performance generally degrades with context perturbations, this is not consistently the case for QASPER due to interactions between abstention and task performance (see Section 5.1).
@@ 5 Results
Baseline task performance in the zero-shot prompting setting is presented in Table 3(a), alongside previously reported SOTA zero-shot performance of LLMs on each dataset. Surprisingly, Flan-T5 achieves the best performance on SQuAD2, PubmedQA, and QASPER, while LLama2 performs best on BioASQ. GPT3.5 achieves the second best performance on PubmedQA, BioASQ, and QASPER; and GPT4o achieves second best performance on SQuAD2, PubmedQA, and BioASQ. Task performance of tested models are close to and sometimes exceed (BioASQ) the reported SOTA on each dataset. Though in this work, we focus on assessing change to task performance under different context settings rather than maximizing model performance.
@@ 5.1 Impact of context perturbations
In Table 4(c), we observe that all models are much more likely to abstain from answering when given random context except on the BioASQ dataset. For SQuAD2 and PubmedQA, random context improves most models’ abstention performance to close to 1. Specifically, random context improves Flan-T5’s ability to abstain from answering on PubmedQA while the no context setting completely fails to do so. For QASPER, the abstention performance on answerable questions in the random context setting is much higher compared to no context, but is still far from 1. Oddly, for boolean questions in BioASQ, random context biases the model towards answering “no” rather than “unanswerable”, and we observe no changes in abstention rate (further analysis in Section 6).
@@
While noisy context has mixed effects on abstention (Table 4(d)), this perturbation does not always translate to negative impacts on task performance. In Table 3(d), task performance increases on QASPER for LLama2 and Vicuna and PubmedQA for Vicuna and GPT4o. For QASPER, unanswerable questions have empty context, so the baseline task performance for unanswerable questions is actually computed with no context; perturbing no context by adding noisy context therefore leads to a trade-off in task performance and abstention performance (some models abstain much more for unanswerable questions and task performance consequently improves).
@@ 5.2 Impact of question type
For extractive questions, Figure 2(a) and (d) show that model abstention varies significantly between given context and no context settings; under no context settings, models can achieve abstention performance close to 1 for a highly context-dependent dataset like QASPER. Abstractive answers show similar patterns. Notably, we find all models are reluctant to abstain from boolean questions for all datasets; this overconfidence is apparent even for QASPER, which is a highly document-specific QA dataset. Surprisingly, Flan-T5 and LLama2 demonstrate near zero abstention performance on boolean questions on any dataset, regardless of whether or not context is provided. [^2] We conduct prompt ablations in Appendix C, showing the lack of abstention on boolean questions is highly sensitive to prompt wording. When a boolean instruction is removed from the prompt, both models are able to abstain much better for boolean questions. Vicuna and Mistral perform slightly better, with some variability across datasets. GPT3.5 and GPT4o demonstrate much higher abstention performance compared to other models, but are still far from 1.
@@
In Figure 2, we examine the abstention performance on unanswerable questions (yellow background) and find that performance varies significantly across datasets and models. For SQuAD2: LLama2, Vicuna, Mistral, and GPT4o abstain less when context is removed, which contradicts intuition. In SQuAD2, unanswerable questions are still related to the provided context; which may explain why when the confusing context is removed, these two models become more likely to answer. GPT3.5 performs consistently regardless of whether context is provided. This consistency suggests that GPT3.5 may rely on internal mechanisms for abstention that are less sensitive to the presence or absence of contextual information. For PubmedQA, unanswerable questions are associated with context that is highly uncertain; Flan-T5 and LLama2 consistently refuse to abstain from answering these questions, which are boolean, while Vicuna follows the “unanswerable” instruction quite well. For QASPER, abstention performance is generally higher across all models compared to SQuAD2 and PubmedQA since the questions are more document-grounded.
@@
GPT4o 22.5 62.0 45.7 38.2
Table 5: Percentage of answerable questions for which the model changes from answering to abstaining when context is removed. Best performances bolded, second best underlined.
5.3 Abstention performance by model
@@
Table 5 shows the percentage of answers for which the model changed from answering to abstaining when context is removed. The expectation for context-dependent questions would be that a model would abstain from answering if no context is provided. GPT4o achieves the largest abstention changes on PubmedQA and BioASQ, while GPT3.5 also abstain well across all four datasets, indicating that they are responsive to context loss. Vicuna, with more instruction-tuning than LLama2, demonstrates relatively strong abstention behavior in these cases compared to LLama2.
@@
Flan-T5 abstains well when context is removed (Table 5). We observe large changes for SQuAD2 and QASPER (comparable or better than GPT4o and GPT3.5), although this behavior does not generalize at all to boolean questions—the change is 0.0 for PubmedQA and BioASQ, which consist only of boolean questions. Different prompting strategies may be necessary to enjoin certain models to abstain when answering boolean questions.
@@ 6 Error Analysis
We perform error analysis on a consistently high performing model in our analysis, GPT3.5, and summarize main reasons why the model may fail to abstain. We sample 20 failure cases each from among answerable questions and unanswerable questions, for both gold context and no context settings. We present our findings by dataset.
@@ 7 Discussion & Conclusion
Our study investigates the impacts of context removal and perturbation on LLM performance on scientific QA. While lack of correct context should result in a model abstaining from answering a question, our results highlight that there are inconsistent patterns of model behavior based on model pretraining paradigms, question types, and the context-dependence of various QA datasets. For example, perturbing given context by replacing with random irrelevant context or adding noisy context would be expected to reduce task performance, but in some cases, the improvements in abstention that result can negate any reductions in task performance and potentially lead to gains (e.g., Vicuna and GPT4o on PubmedQA, LLama2 and Vicuna on QASPER). Additionally, we find that abstention varies greatly by question type, with all models in our experiments struggling to abstain on boolean questions.
@@
Future work for enhancing models’ abstention ability could investigate the impact of (i) Different prompting strategies: since model abstention ability is sensitive to constrained prompts (such as boolean instructions) as shown in Section 5.2, how to select good prompting strategies to produce the best trade-off between task performance and abstention performance remains an interesting problem; (ii) Alternate model architectures: smaller models such as Flan-T5-XL with encoder-decoder architectures performed comparably to larger decoder-only models such as ChatGPT. Further exploration of encoder-decoder architectures or introducing an auxiliary module to foster understanding of context may be helpful; (iii) Other context perturbations: we show initial results that providing noisy context can counter-intuitively improve task performance in some datasets due to interactions with unanswerable questions. How this interacts with retrieval errors or malicious injected context that occur in open-domain QA is a direction that could be explored in future work.
@@
Our results have direct implications for dataset curators and model developers. Benchmark QA datasets with mixtures of unanswerable and answerable questions were designed to facilitate assessment of abstention ability, yet our experiments show conflation and a lack of clear assessment of either abstention ability or performance accuracy. In other words, while unanswerable questions were motivated by the need to measure abstention, when aggregated with other questions during task performance evaluation, this distinction is obscured. Additionally, unanswerable questions in different datasets measure different phenomena, e.g., in SQuAD2, they measure model sensitivity to irrelevant context passages, while in QASPER, they indicate that a question is unanswerable based on the given document, neither of which clearly map to the notion of abstaining under insufficient information. Given the importance of assessing model abstention capabilities, separating task and abstention assessment, coupled with changes in dataset construction, is needed to better align model performance on these tasks with human expectations.
@@
On the other hand, our results also hold implications for downstream builders/users of interactive systems who rely on LLMs as question-answering tools. We show that in some cases (e.g., boolean questions), LLMs exhibit little to no ability to abstain, even for highly context-dependent questions such as those in QASPER. System designers should be cautious of these limitations and when/how to infer model uncertainty and communicate this uncertainty to users. For communicating abstention capabilities to users, confidence scores or similar indicators that reflect model certainty, accompanying explanations, or interpretation guidance can empower users to make more informed decisions.
@@
Due to the budget constraints, we could not conduct experiments with every single model, dataset, or context perturbation method from existing work, though our framework can be applied to other datasets and models. We will make our framework publicly available such that new models and datasets, as well as context perturbations could be added. Other factors affecting abstention such as in-domain pretraining and in-context learning could be incorporated in the future.
Acknowledgements
This research was supported in part by the UW iSchool Strategic Research Fund.